# Unbiased Regression-Adjusted Estimation of Average Treatment Effects in Randomized Controlled Trials

### v1

@@ B.2 Omitted Proofs of Section˜4
$\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}:=\arg\min_{{\bm{b}}\in\mathbb{R}^{d}}\left\|{\bm{X}}{\bm{b}}-\widetilde{{\bm{y}}}^{(-i)}\right\|_{2}^{2}+\lambda\|{\bm{b}}\|_{2}^{2}=({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\widetilde{{\bm{y}}}^{(-i)}.$
Unlike $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$, this uses all $n$ rows of $\bm{X}$} rather than dropping unit $i$. We show that $\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}\to\boldsymbol{\beta}^{*}_{\lambda}$ in probability (with probability taken over the random assignment).
Algebraic manipulation yields $\widetilde{{\bm{y}}}^{(-i)}=\frac{\widetilde{{\boldsymbol{\mu}}}}{n}+\frac{{\bm{v}}^{(-i)}{\bm{z}}\widetilde{{\bm{t}}}^{(-i)}}{n}$ and ${\mathbb{E}}\big[\widetilde{{\bm{y}}}^{(-i)}\big]=\frac{\widetilde{{\boldsymbol{\mu}}}}{n}$. We denote the deviation with
$\Delta^{(-i)}\;:=\;\widetilde{{\bm{y}}}^{(-i)}-\frac{\widetilde{{\boldsymbol{\mu}}}}{n}.$
Because assignment is complete random assignment (sampling without replacement), the dependence across units is weak: covariances between $d_{j}$ and $d_{k}$ are of order $1/n$ for $j\neq k$. Using this standard finite-population sampling algebra and the bounded fourth-moment assumptions in Assumptions˜1 and 2, one obtains
$\big\|{\bm{X}}^{\top}\Delta^{(-i)}\big\|_{2}=O(\sqrt{n}),$
uniformly in $i$. Intuitively, $\Delta^{(-i)}$ is a mean-zero perturbation with components of order $1$, and summing ${\bm{x}}_{j}\Delta^{(-i)}_{j}$ accumulates fluctuations of order $\sqrt{n}$ in Euclidean norm.
We may now write
${\bm{X}}^{\top}\widetilde{{\bm{y}}}^{(-i)}={\bm{X}}^{\top}\frac{\widetilde{{\boldsymbol{\mu}}}}{n}+{\bm{X}}^{\top}\Delta^{(-i)}.$
Premultiplying by $({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}$ gives
$\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}=({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\frac{\widetilde{{\boldsymbol{\mu}}}}{n}\;+\;({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\Delta^{(-i)}.$
The first term converges to $\boldsymbol{\beta}^{*}_{\lambda}$ by Step 1. For the second term, we use $\|({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}\|_{2}=O(n^{-1})$ and $\|{\bm{X}}^{\top}\Delta^{(-i)}\|_{2}=O(\sqrt{n})$, which implies
$\left\|({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\Delta^{(-i)}\right\|_{2}=O(n^{-1})\cdot O(\sqrt{n})=O(n^{-1/2})\;\to\;0.$
Hence
$\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}\;\xrightarrow{p}\;\boldsymbol{\beta}^{*}_{\lambda},\qquad\text{for each }i.$
Step 3: Leave-one-out versus full-sample ridge.
We now compare $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$, which regresses $\widetilde{{\bm{y}}}^{(-i)}_{-i}$ on ${\bm{X}}_{-i}$, to $\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}$, which regresses $\widetilde{{\bm{y}}}^{(-i)}$ on $\bm{X}$}. The standard leave-one-out identity for ridge regression (cf. Lemma˜B.1) yields
@@
$\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}=\frac{({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}\left(\widetilde{y}^{(-i)}_{i}-{\bm{x}}_{i}^{\top}\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}\right)}{1-h_{\lambda ii}},$ (39)
@@
$h_{\lambda ii}={\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}$
is the ridge leverage score of unit $i$. By Assumption˜2, $\|{\bm{x}}_{i}\|_{2}$ is uniformly bounded. Using (38),
$h_{\lambda ii}\;\leq\;\|{\bm{x}}_{i}\|_{2}^{2}\,\big\|({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}\big\|_{2}=O(n^{-1}),$
so $1-h_{\lambda ii}\to 1$ uniformly in $i$. Moreover, $\|({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}\|_{2}=O(n^{-1})$, and $\big|\widetilde{y}^{(-i)}_{i}-{\bm{x}}_{i}^{\top}\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}\big|=O(1)$, because $\widetilde{y}^{(-i)}_{i}$ is a reweighting of $y_{i}^{(1)}$ or $y_{i}^{(0)}$ with coefficients of order $1$, and $\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ is bounded by the conclusion in Step 2. Thus the right-hand side of (39) is $O(n^{-1})$, uniformly in $i$, and therefore
@@
$\big\|\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\big\|_{2}\;\to\;0.$
Finally, Step 2 shows $\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}\xrightarrow{p}\boldsymbol{\beta}^{*}_{\lambda}$ and Step 3 shows $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widetilde{\boldsymbol{\beta}}^{(-i)}_{\lambda}\to\boldsymbol{0}$. Combining these, for each $i$,
$\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\;\xrightarrow{p}\;\boldsymbol{\beta}^{*}_{\lambda}.$
@@
Recall that
@@
$\widehat{\tau}_{\textup{LDM}}=\frac{1}{n_{T}}\sum_{i:d_{i}=1}\big(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\big)\;-\;\frac{1}{n_{C}}\sum_{i:d_{i}=0}\big(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\big)$,
where $d_{i}\in\{0,1\}$ is the treatment indicator, $n_{T}=\sum_{i}d_{i}$, $n_{C}=n-n_{T}$, and $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ is the leave-one-out ridge coefficient for unit $i$.
By Lemma˜6,
$\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\;\xrightarrow{p}\;\boldsymbol{\beta}^{*}_{\lambda}\quad\text{for each }i.$
Define the infeasible “oracle” regression-adjusted difference-in-means estimator that uses $\boldsymbol{\beta}^{*}_{\lambda}$ in place of $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$:

### v2

@@ B.2 Omitted Proofs of Section˜4
$\boldsymbol{\beta}_{\lambda}$ $=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\left\|{\bm{X}}{\bm{b}}-\widetilde{{\boldsymbol{\mu}}}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2}.$
@@
$\widehat{\boldsymbol{\beta}}_{\lambda}^{(i)}$ $=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\left\|{\bm{X}}{\bm{b}}-\widetilde{{\bm{y}}}^{(-i)}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2}.$
$\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$ $=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\left\|{\bm{X}}_{-i}{\bm{b}}-\widetilde{{\bm{y}}}_{-i}^{(-i)}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2}.$
Define the oracle difference-in-means estimator as
$\widehat{\tau}_{\textup{LDM}}^{*}=\sum_{i=1}^{n}v_{i}z_{i}\bigl(y_{i}-\eta^{-1}{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}\bigr).$
Note that LOORA-DM estimator is
$\widehat{\tau}_{\textup{LDM}}=\sum_{i=1}^{n}v_{i}z_{i}\bigl(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\bigr)$.
We define
$\frac{\sqrt{n}\bigl(\widehat{\tau}_{\textup{LDM}}-\tau\bigr)}{\sqrt{{\rm Var}(\sqrt{n}(\widehat{\tau}_{\textup{LDM}}^{*}-\tau))}}=A_{n}+B_{n},~A_{n}=\frac{\sqrt{n}\bigl(\widehat{\tau}_{\textup{LDM}}^{*}-\tau\bigr)}{\sqrt{{\rm Var}(\sqrt{n}(\widehat{\tau}_{\textup{LDM}}^{*}-\tau))}},$~
$B_{n}=\frac{\sqrt{n}\bigl(\widehat{\tau}_{\textup{LDM}}-\widehat{\tau}_{\textup{LDM}}^{*}\bigr)}{\sqrt{{\rm Var}(\sqrt{n}(\widehat{\tau}_{\textup{LDM}}^{*}-\tau))}}.$
Note that
$\sqrt{{\rm Var}(\sqrt{n}(\widehat{\tau}_{\textup{LDM}}^{*}-\tau))}=\frac{\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{\lambda}\allowbreak-\allowbreak(\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}})\right\|_{2}}{\sqrt{n-1}}.$
Step 1: Negligibility of the plug-in error ($B_{n}=o_{p}(1)$).
We have
$\widehat{\tau}_{\textup{LDM}}-\widehat{\tau}_{\textup{LDM}}^{*}=\sum_{i=1}^{n}v_{i}z_{i}\,{\bm{x}}_{i}^{\top}(\eta^{-1}\boldsymbol{\beta}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}).$
Add and subtract $\widehat{\boldsymbol{\beta}}_{\lambda}$:
$\left|\widehat{\tau}_{\textup{LDM}}-\widehat{\tau}_{\textup{LDM}}^{*}\right|\leq\underbrace{\left|\sum_{i=1}^{n}v_{i}z_{i}\,{\bm{x}}_{i}^{\top}(\eta^{-1}\boldsymbol{\beta}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda}^{(i)})\right|}_{T_{1}}+\underbrace{\left|\sum_{i=1}^{n}v_{i}z_{i}\,{\bm{x}}_{i}^{\top}(\widehat{\boldsymbol{\beta}}_{\lambda}^{(i)}-\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})\right|}_{T_{2}}.$
Note that by (23),
$\eta~\widetilde{{\bm{y}}}^{(-i)}=\widetilde{{\boldsymbol{\mu}}}+\eta\frac{{\bm{v}}^{(-i)}{\bm{z}}\widetilde{{\bm{t}}}^{(-i)}}{n}.$
Therefore
$\eta^{-1}\boldsymbol{\beta}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda}^{(i)}=-({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\frac{{\bm{v}}^{(-i)}{\bm{z}}\widetilde{{\bm{t}}}^{(-i)}}{n}.$
Let $\widetilde{\boldsymbol{\beta}}$ be a vector such that
$\eta^{-1}\boldsymbol{\beta}_{\lambda}-\widetilde{\boldsymbol{\beta}}_{\lambda}=-({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\frac{{\bm{v}}{\bm{z}}\widetilde{{\bm{t}}}}{n},$
@@
$\widetilde{{\bm{t}}}=n_{C}(n_{C}-1){\bm{y}}^{(1)}-n_{T}(n_{T}-1){\bm{y}}^{(0)}.$
Then one can easily show that
$\left\|\widehat{\boldsymbol{\beta}}_{\lambda}^{(i)}-\widetilde{\boldsymbol{\beta}}_{\lambda}\right\|_{2}=O(n^{-1}).$
Now by triangle inequality and Cauchy–Schwarz inequality,
$T_{1}$ $\leq\left|\sum_{i=1}^{n}v_{i}z_{i}\,{\bm{x}}_{i}^{\top}(\eta^{-1}\boldsymbol{\beta}_{\lambda}-\widetilde{\boldsymbol{\beta}}_{\lambda})\right|+\left|\sum_{i=1}^{n}v_{i}z_{i}\,{\bm{x}}_{i}^{\top}(\widetilde{\boldsymbol{\beta}}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda}^{(i)})\right|$
$=\left|\sum_{i=1}^{n}v_{i}z_{i}\,{\bm{x}}_{i}^{\top}(\eta^{-1}\boldsymbol{\beta}_{\lambda}-\widetilde{\boldsymbol{\beta}}_{\lambda})\right|+O(n^{-1})$
$=\left|({\bm{X}}^{\top}({\bm{v}}{\bm{z}}))^{\top}(\eta^{-1}\boldsymbol{\beta}_{\lambda}-\widetilde{\boldsymbol{\beta}}_{\lambda})\right|+O(n^{-1})$
$\leq\left\|{\bm{X}}^{\top}({\bm{v}}{\bm{z}})\right\|_{2}\left\|\eta^{-1}\boldsymbol{\beta}_{\lambda}-\widetilde{\boldsymbol{\beta}}_{\lambda}\right\|_{2}+O(n^{-1})$
$\leq\left\|{\bm{X}}^{\top}({\bm{v}}{\bm{z}})\right\|_{2}\left\|({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}\right\|_{2}\left\|{\bm{X}}^{\top}\frac{{\bm{v}}{\bm{z}}\widetilde{{\bm{t}}}}{n}\right\|_{2}+O(n^{-1})=O_{p}(n^{-1}),$
where the first equality holds because $v_{i}=O(n^{-1})$ for all $i\in[n]$ and the last equality holds due to Lemma˜9 and noting that
$v_{i}z_{i}=\frac{d_{i}}{n_{T}}-\frac{1-d_{i}}{n_{C}}=\Bigl(\frac{1}{n_{T}}+\frac{1}{n_{C}}\Bigr)(d_{i}-p_{T}).$
Now note that by Lemma˜6 and because of Assumptions˜1, 2 and 3
$\left\|\widehat{\boldsymbol{\beta}}_{\lambda}^{(i)}-\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\right\|_{2}=O(n^{-1}).$ (27)
Therefore by triangle inequality and Cauchy–Schwarz inequality,
$T_{2}\leq\sum_{i=1}^{n}\left\|v_{i}z_{i}\,{\bm{x}}_{i}^{\top}\right\|_{2}\left\|\widehat{\boldsymbol{\beta}}_{\lambda}^{(i)}-\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\right\|_{2}=O(n^{-1}),$
where the equality holds because $v_{i}=O(n^{-1})$ for all $i\in[n]$. Therefore $\left|\widehat{\tau}_{\textup{LDM}}-\widehat{\tau}_{\textup{LDM}}^{*}\right|=O_{p}(n^{-1})$ and due to the assumption that $\left\|\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}}-({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}\right\|_{2}\to\infty$, $B_{n}=o_{p}(1)$.
Step 2: Oracle CLT via Hoeffding’s combinatorial CLT.
Let
$u_{i}^{(1)}=y_{i}^{(1)}-\eta^{-1}{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda},~~\text{and}~~u_{i}^{(0)}=y_{i}^{(0)}-\eta^{-1}{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}$
Therefore we have
$\widehat{\tau}_{\textup{LDM}}^{*}$ $=\frac{1}{n_{T}}\sum_{i\in[n]:d_{i}=1}u_{i}^{(1)}-\frac{1}{n_{C}}\sum_{i\in[n]:d_{i}=0}u_{i}^{(0)}$
$=-\frac{1}{n_{C}}\sum_{i=1}^{n}u_{i}^{(0)}+\sum_{i\in[n]:d_{i}=1}\left(\frac{1}{n_{T}}u_{i}^{(1)}+\frac{1}{n_{C}}u_{i}^{(0)}\right)$
@@
$=-\frac{1}{n_{C}}\sum_{i=1}^{n}u_{i}^{(0)}+\frac{1}{\sqrt{n_{T}n_{C}}}\sum_{i\in[n]:d_{i}=1}\left(\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}\right).$
The first term on the right-hand-side of the above is determinstic and the only sourse of randomness is the second term.
Suppose for each natural number $n$, we have $2n$ numbers $a_{ni},c_{ni}$. For a permutation $\pi(1),\ldots,\pi(n)$ of $[n]$, define the following random variable.
$S_{n}=\sum_{i=1}^{n}a_{ni}c_{n\pi(i)}.$
Hoeffding (1951) shows that
$\frac{S_{n}-{\mathbb{E}}(S_{n})}{\sqrt{{\rm Var}(S_{n})}}\;\xrightarrow{d}\;\mathcal{N}(0,1)$
if
$\lim_{n\to\infty}n\cdot\frac{\max_{i\in[n]}(a_{ni}-\overline{a}_{n})^{2}}{\sum_{i=1}^{n}(a_{ni}-\overline{a}_{n})^{2}}\cdot\frac{\max_{i\in[n]}(c_{ni}-\overline{c}_{n})^{2}}{\sum_{i=1}^{n}(c_{ni}-\overline{c}_{n})^{2}}=0,$
where
$\overline{a}_{n}=\frac{1}{n}\sum_{i=1}^{n}a_{ni},~~\text{and}~~\overline{c}_{n}=\frac{1}{n}\sum_{i=1}^{n}c_{ni}.$
We set $a_{ni}=(\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})$ and $c_{ni}=\frac{1}{\sqrt{n_{T}n_{C}}}\mathbbm{1}\{i\leq n_{T}\}$, where $\mathbbm{1}\{i\leq n_{T}\}$ is an indicator function for $i\leq n_{T}$. Assumptions˜1, 2 and 3 imply that each $a_{ni}$ is bounded. Therefore each $a_{ni}-\overline{a}_{n}$ is bounded. Let $C$ be a constant such that $|a_{ni}-\overline{a}_{n}|\leq C$ for all $n$ and $i$. Then
$\frac{\max_{i\in[n]}(a_{ni}-\overline{a}_{n})^{2}}{\sum_{i=1}^{n}(a_{ni}-\overline{a}_{n})^{2}}$ $\leq\frac{C^{2}}{\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{\lambda}\allowbreak-\allowbreak(\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}})\right\|_{2}}.$
Now note that $\overline{c}_{n}=\frac{1}{\sqrt{n_{T}n_{C}}}\frac{n_{T}}{n}$ and therefore each $c_{ni}-\overline{c}_{n}$ is either equal to $\frac{1}{\sqrt{n_{T}n_{C}}}\frac{n_{C}}{n}$ or $\frac{1}{\sqrt{n_{T}n_{C}}}\frac{n_{T}}{n}$. Therefore we have
$\frac{\max_{i\in[n]}(c_{ni}-\overline{c}_{n})^{2}}{\sum_{i=1}^{n}(c_{ni}-\overline{c}_{n})^{2}}$ $=\frac{\max\{\left(\frac{n_{C}}{n}\right)^{2},\left(\frac{n_{T}}{n}\right)^{2}\}}{n_{T}\left(\frac{n_{C}}{n}\right)^{2}+n_{C}\left(\frac{n_{T}}{n}\right)^{2}}=\frac{\max\{n_{C}^{2},n_{T}^{2}\}}{n_{C}n_{T}n}=\frac{1}{n}\max\{\frac{n_{C}}{n_{T}},\frac{n_{T}}{n_{C}}\}$
$\leq\frac{1}{n}\frac{1-m}{m}.$
where the inequality follows from Assumption˜2. Thus
$n\cdot\frac{\max_{i\in[n]}(a_{ni}-\overline{a}_{n})^{2}}{\sum_{i=1}^{n}(a_{ni}-\overline{a}_{n})^{2}}\cdot\frac{\max_{i\in[n]}(c_{ni}-\overline{c}_{n})^{2}}{\sum_{i=1}^{n}(c_{ni}-\overline{c}_{n})^{2}}\leq\frac{C^{2}}{\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{\lambda}\allowbreak-\allowbreak(\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}})\right\|_{2}}\frac{1-m}{m}.$
Therefore because of the assumption that $\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{\lambda}\allowbreak-\allowbreak(\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}})\right\|_{2}\to\infty$, the limit of the above is zero. Therefore by Hoeffding (1951),
$\frac{\sqrt{n}\bigl(\widehat{\tau}_{\textup{LDM}}^{*}-\tau\bigr)}{\sqrt{{\rm Var}(\sqrt{n}(\widehat{\tau}_{\textup{LDM}}^{*}-\tau))}}\;\xrightarrow{d}\;\mathcal{N}(0,1).$
Moreover note that
$\sqrt{{\rm Var}(\sqrt{n}(\widehat{\tau}_{\textup{LDM}}^{*}-\tau))}=\frac{\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{\lambda}\allowbreak-\allowbreak(\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}})\right\|_{2}}{\sqrt{n-1}}.$
Step 3: Conclusion.
Since $A_{n}\xrightarrow{d}\mathcal{N}(0,1)$ and $B_{n}=o_{p}(1)$, Slutsky’s theorem gives
$\frac{\sqrt{n}\bigl(\widehat{\tau}_{\textup{LDM}}-\tau\bigr)}{\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{\lambda}\allowbreak-\allowbreak(\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}})\right\|_{2}/\sqrt{n}}\;\xrightarrow{d}\;\mathcal{N}(0,1),$
where we also use the fact that $\lim_{n\to\infty}\frac{n}{n-1}=1$. ∎
@@
We define
$\boldsymbol{\beta}_{\lambda}$ $=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\left\|{\bm{X}}{\bm{b}}-\widetilde{{\boldsymbol{\mu}}}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2},$
$\widehat{\boldsymbol{\beta}}_{\lambda}^{(i)}$ $=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\left\|{\bm{X}}{\bm{b}}-\widetilde{{\bm{y}}}^{(-i)}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2},$
$\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$ $=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\left\|{\bm{X}}_{-i}{\bm{b}}-\widetilde{{\bm{y}}}_{-i}^{(-i)}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2}.$
and
$\widetilde{V}_{\textup{LDM}}=\frac{n-1}{n_{C}(n_{C}-1)}\sum_{i\in[n]:d_{i}=0}(\widetilde{s}_{i}^{(0)})^{2}+\frac{n-1}{n_{T}(n_{T}-1)}\sum_{i\in[n]:d_{i}=1}(\widetilde{s}_{i}^{(1)})^{2},$
where
$\widetilde{s}_{i}^{(0)}$ $=y_{i}^{(0)}-\eta^{-1}{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}-\frac{1}{n_{C}}\sum_{j\in[n]:d_{j}=0}y_{j}^{(0)}-\eta^{-1}{\bm{x}}_{j}^{\top}\boldsymbol{\beta}_{\lambda},$
$\widetilde{s}_{i}^{(1)}$ $=y_{i}^{(1)}-\eta^{-1}{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}-\frac{1}{n_{T}}\sum_{j\in[n]:d_{j}=1}y_{j}^{(1)}-\eta^{-1}{\bm{x}}_{j}^{\top}\boldsymbol{\beta}_{\lambda}.$
We show
@@
$\widehat{V}_{\textup{LDM}}-\left(\frac{1}{n}\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{\lambda}-(\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}})\right\|_{2}^{2}+\frac{1}{n}\sum_{i=1}^{n}(\tau_{i}-\tau)^{2}\right)\xrightarrow{p}0$
by separating the left-hand-side to two terms: $\widehat{V}_{\textup{LDM}}-\widetilde{V}_{\textup{LDM}}$ and